Use the definition of limit to prove that both of the sequences { 1 √ n √ } and {(−1)n n } converges to 0.


1
Expert's answer
2021-01-26T04:18:28-0500

Let "a_{n}=\frac{1}{\sqrt{n}}" . Let "\epsilon>0" is given. Then there exist some N>0 satisfying "N>\frac{1}{\epsilon^{2}}".( Archimedean property of real numbers)


Therefore "\forall" "n\geq N",


"|a_{n}-0|=|\frac{1}{\sqrt{n}}|=\frac{1}{\sqrt{n}}\leq \frac{1}{\sqrt{N}}<\epsilon" .


Hence from the definition of limits "a_{n} \rightarrow 0."


Let "b_{n}= \frac{(-1)^{n}}{n}." Then "\forall n\geq N",


"|b_{n}-0|=|\frac{(-1)^{n}}{n}|=\frac{1}{n}\leq \frac{1}{N}<\epsilon" .


Hence from the definition of limits "b_{n}\rightarrow 0."